# Lizhang Chen*†‡ Jonathan Li*† Kaizhao Liang Baiyu Su

### v1

@@ 1 Introduction
Optimization algorithms lie at the core of modern deep learning, shaping not only convergence speed but also training stability and generalization ability across domains such as natural language processing and computer vision. As models and datasets scale, traditional methods such as stochastic gradient descent (SGD) and SGD with momentum (SMDH, 13) encounter limitations, including slow convergence in non-convex landscapes, sensitivity to learning rate schedules, and poor robustness to sparse or noisy gradients (SM, 20; ZMB+, 25). In response, a wide range of alternatives have emerged, including adaptive gradient methods (DHS, 11; KB, 15), approximate second-order approaches (MG, 15; GKS, 18; YGS+, 21; LLH+, 24; NCLL, 24; WHML, 25), and specialized algorithms for extreme training regimes (LLCL, 24; LYL, 24; XZL+, 24; HZJ+, 25; ZCL+, 25).
Among these advances, decoupled weight decay (LH, 19) has proven especially influential. In its general form, decoupled weight decay augments any optimizer update $\mathbf{u}_{t}$ with a decay term applied directly to the parameters, i.e.
@@
This technique improves training stability and generalization by preventing the adaptive learning rates from interfering with regularization, as exemplified by the success of AdamW in large model training (BMR+, 20; DBK+, 21; TMS+, 23) and the subsequent development of state-of-the-art optimizers such as Lion (CLH+, 23), Lion-$\mathcal{K}$ (CLLL, 24), and Muon (JJB+, 24; LSY+, 25).
However, decoupled weight decay remains agnostic to the directional alignment between the optimizer update and the parameters, which may hurt performance when they conflict. Intuitively, when the update $\mathbf{u}_{t}$ and parameters $\mathbf{x}_{t}$ point in the same direction for a given dimension, weight decay acts as a regularizer that improves stability; however, when their directions differ, applying decay actively resists beneficial movement toward the optimum. Furthermore, decoupled weight decay has been shown to implicitly impose regularization terms on the objective function (CLLL, 24; XL, 24), which corresponds to parameter norm constraints for AdamW, Lion, and Muon.
@@
$\bullet$ We use Lyapunov analysis to show that standard optimizers (SGD(M), Lion-$\mathcal{K}$, Adam) with cautious weight decay are asymptotically stable and unbiased, in the sense that they optimize the original loss rather than a regularized surrogate. The regularization effect of cautious weight decay instead becomes a bilevel objective of finding locally Pareto-optimal points within the stationary manifold (Figure 2). Furthermore, we show that discrete-time Adam with cautious weight decay attains a standard convergence rate in the smooth nonconvex setting.
$\bullet$ In language modeling (OWS+, 25; KFP+, 25) and ImageNet classification (DDS+, 09), we observe that cautious weight decay generally accelerates convergence and lowers final validation loss for AdamW, Lion, and Muon (e.g., Figure 1). These improvements translate into higher zero-shot accuracy on standard benchmarks from 338M to 2B parameters and across architectures without retuning baseline settings ($\approx$20,000 NVIDIA H100 HBM3-80GB GPU hours for all experiments).
@@ 2.1 Decoupled weight decay
where $\beta_{1}$ and $\beta_{2}$ are momentum coefficients and $\nabla\mathcal{K}$ is a subgradient of $\mathcal{K}$. Examples include Lion when $\mathcal{K}=\left\lVert\cdot\right\rVert_{1}$ and Muon when $\mathcal{K}=\left\lVert\cdot\right\rVert_{\textup{tr}}$, where $\left\lVert\cdot\right\rVert_{\textup{tr}}$ denotes the trace norm when the parameters are treated as a matrix.
@@ 2.2 Implicit regularization effects of weight decay
This suggests that Lion-$\mathcal{K}$ optimizes the regularized objective $\widehat{f}$, an observation made by CLLL (24). In the special cases of Lion and Muon, $\mathcal{K}^{*}$ is the 0-$\infty$ indicator function of a dual norm ball, corresponding to the constrained optimization problems
@@
A similar analysis for AdamW suggests that it solves the box-constrained problem of minimizing $f(\mathbf{x})$ such that $\left\lVert\mathbf{x}\right\rVert_{\infty}\leq\frac{1}{\lambda}$, but convergence cannot be established due to the lack of a Lyapunov function. For more discussion, see Appendix C and XL (24).
@@ 3.1 Convergence to the stationary manifold
where $(\cdot)^{+}:=\max(0,\cdot)$. LaSalle’s invariance principle (LaS, 60) states that the accumulation points of any trajectory lie within the union of trajectories $\mathbf{z}_{t}$ that satisfy $\tfrac{\textup{d}}{\textup{d}t}\mathcal{H}(\mathbf{z}_{t})=0$ for all $t\geq 0$. Consequently, we conclude that SGD with cautious weight decay produces trajectories that approach the stationary set $\{\mathbf{x}\mid\nabla f(\mathbf{x})=\mathbf{0}\}$ of the original loss. This holds because cautious weight decay is applied only in a secondary fashion and is automatically deactivated whenever it conflicts with the main objective, thereby ensuring that the loss landscape remains unbiased.
@@ 3.2 Sliding mode dynamics
Due to the discontinuity of $f_{t}$, the solution may not be well defined in the classical or Carathéodory sense, especially across switching surfaces. We therefore interpret solutions in the Filippov sense (Fil, 88), where a discontinuous ODE is formally a differential inclusion that specifies that $\dot{\mathbf{z}}_{t}$ belongs to the closed convex envelope of the discontinuous vector field, i.e.
@@
Figure 3: Toy objectives and trajectories. Left: $f(x,y)=((y-3)^{2}-(x-3)^{2}-1)^{2}$. Right: $f(x,y)=(y-3-(x-3)^{2})^{2}$. We compare Adam, AdamW, and Adam + CWD; AdamW and CWD use the same weight decay $\lambda$, and all other hyperparameters $(\eta,\beta_{1},\beta_{2},\epsilon)$ are identical. For both objectives, Adam converges to a generic point on the minimizer manifold, whereas AdamW converges to a solution of the box-constrained problem $\min_{x,y}f(x,y)$ subject to $\max\{x,y\}\leq\frac{1}{\lambda}$. In contrast, Adam + CWD converges to the Pareto front of the minimizer manifold.
4 Discrete-Time Analysis
Algorithm 2 Adam with cautious weight decay
3:  repeat
7:   $\widehat{\mathbf{m}}_{t}\leftarrow(1-\beta_{1}^{t})^{-1}\mathbf{m}_{t}$
8:   $\widehat{\mathbf{v}}_{t}\leftarrow(1-\beta_{2}^{t})^{-1}\mathbf{v}_{t}$
[…]
10:   $t\leftarrow t+1$
11:  until stopping criterion is met
12:  return optimized parameters $\mathbf{x}_{t}$
Leveraging the Lyapunov functions in Section 3, it is possible to extend our analysis to the discrete-time dynamics of various optimizers with cautious weight decay. In this section, we use Adam with cautious weight decay (Algorithm 2) as an example, showing that in the smooth nonconvex setting, Algorithm 2 achieves a standard convergence rate of $O(T^{-\frac{1}{2}})$ on the squared gradient norm and an additional stationarity measure.
We make the following assumptions, which are mild and often used in the analysis of stochastic gradient algorithms (GL, 13; BB, 21; DBBU, 22; ACD+, 23).
Assumption 1.
$f$ is coercive and $L$-smooth. This implies that $f$ attains a minimum value, which we denote as $f^{\star}$, and that the iterates of Algorithm 2 are bounded.
Assumption 2 (Bounded variance).
The stochastic gradient $\mathbf{g}_{t}$ satisfies
$\mathbb{E}[\mathbf{g}_{t}]=\nabla f(\mathbf{x}_{t})\quad\text{and}\quad\textup{Var}(\mathbf{g}_{t})=\mathbb{E}\left[\left\lVert\mathbf{g}_{t}-\nabla f(\mathbf{x}_{t})\right\rVert_{2}^{2}\right]\leq\frac{\sigma^{2}}{n_{\mathrm{batch}}},$
where $\sigma$ is a constant and $n_{\mathrm{batch}}$ denotes the batch size.
Theorem 1.
Under Assumptions 1 and 2, let $0\leq\beta_{1}\leq\beta_{2}<1$, $\lambda\geq 0$, $\epsilon>0$, and $\eta_{t}=\eta>0$, and suppose $\mathbf{x}_{t}$ is updated using Algorithm 2. Then for all $T\in\mathbb{N}$,
where $K_{1}$, $K_{2}$, $K_{3}$, and $K_{4}$ are constants.
Proof sketch.
We follow the standard approach of first proving a descent lemma. The full proof is deferred to Appendix E. ∎
Remark 1.
Remark 2.
In the setting of Theorem 1, let $T\in\mathbb{N}$, $\eta=\Theta\left(\frac{1}{\sqrt{T}}\right)$, and $n_{\mathrm{batch}}=\Theta(T)$. Then
$\frac{1}{T}\sum_{t\in[T]}\mathbb{E}\left[\left\lVert\nabla f(\mathbf{x}_{t})\right\rVert_{2}^{2}+\lambda\left\lVert(\nabla f(\mathbf{x}_{t})\mathbf{x}_{t})^{+}\right\rVert_{1}\right]=O\left(\frac{1}{\sqrt{T}}\right).$
@@ 5 Experiments
Overview. We evaluate CWD against three standard optimizers—AdamW, Lion, and Muon—on autoregressive language modeling and ImageNet classification. For Transformer models with similar architecture to Gemma (KFP+, 25) with 338M, 986M, and 2B parameters in the Simply (LHY+, 25) codebase, we follow the Chinchilla compute-optimal scaling rule—20 tokens per parameter (TPP) (HBM+, 22) and train on C4 (RSR+, 20). For each size, we grid-search batch size, learning rate, weight decay, warmup ratio, and optimizer-specific hyperparameters for the baselines (AdamW, Lion, Muon); we then reuse the selected baseline settings for CWD without retuning (details in Appendix F). Under matched settings, CWD lowers final validation loss and improves zero-shot accuracy. On the OLMo codebase (OWS+, 25), we further study an over-training regime—OLMo-1B trained on 100B tokens (100 TPP) from Dolma (SKB+, 24). Under matched settings, CWD lowers final validation loss and improves zero-shot accuracy (Table 4). We also observe similar gains on ImageNet (DDS+, 09) across ViT (DBK+, 21) and ResNet (HZRS, 16).
@@
Ablations on masking. Table 2 tests whether the benefits arise from the amount of decay applied or from CWD’s structure. Replacing our mask with a time-matched Bernoulli “random mask” substantially degrades performance (e.g., $2.56\!\to\!2.82$ for AdamW, $2.42\!\to\!2.73$ for Muon), showing that simply reducing the frequency of decay is insufficient. Substituting the indicator with the gradient-based $\mathbb{I}(\mathbf{g}\mathbf{x}\geq 0)$ also underperforms. Finally, $\lambda=0$ remains worse than tuned decay, illustrating that explicit regularization is helpful and CWD leverages it more effectively.
@@ 6 Related Work
Weight decay. Weight decay originated as an $\ell_{2}$ penalty for ill-posed problems and ridge regression (Tik, 63; HK, 70) and was introduced to neural networks as a generalization tool to mitigate overfitting (HP, 88; WRH, 90; KH, 91). LH (19) showed that, for adaptive methods, weight decay and $\ell_{2}$ are not equivalent, motivating the decoupled formulation in AdamW; subsequent work established decoupled decay as a standard feature of modern optimizers (CLH+, 23; CLLL, 24; LSY+, 25). Recent analyses suggest that in contemporary networks, weight decay functions more as a training accelerator and stabilizer than as explicit regularization (KSH, 17; HBM+, 22; PC, 23; DAVF, 24). Interactions with normalization layers and learning rate schedules have also been clarified (Def, 25), and architectural designs can obviate explicit decay (LHSG, 25).
@@
Weight decay variants. Various efforts have been made to develop different adaptive variants of weight decay. For example, XXZ+ (23) found that weight decay can lead to large gradient norms at the final phase of training and proposed Scheduled Weight Decay (SWD) to dynamically adjust weight decay strength based on gradient norms. KMJ (24) investigate how weight decay affects individual neuron updates, revealing rotational equilibrium states that balance learning across layers and neurons. GSA (23) introduce adaptive weight decay that automatically tunes the hyperparameter during training based on classification and regularization loss gradients, achieving significant improvements in adversarial robustness.
Masked or conditional updates.
@@
Several works have explored the sign-based conditioning of optimizer updates. RB (93) introduced Rprop, which adjusted step sizes based on current gradient and past gradient sign agreement. LCLL (24) propose the cautious optimizer, which restricts updates to dimensions where the proposed update and current gradient share the same sign. WLX+ (24) apply a similar mask to Adam to improve robustness in online learning. Our work is the first to integrate the intuition behind these approaches with decoupled weight decay, showing that selective decay is not merely compatible with adaptive methods but can enhance them significantly.
@@
Constrained and bilevel optimization. Decoupled weight decay can be interpreted through the lens of Frank–Wolfe algorithms for constrained optimization (FW, 56; Jag, 13; SW, 25; PXA+, 25). This connection suggests that optimizers with decoupled weight decay implicitly solve constrained optimization problems, which was shown to be the case for Lion (CLLL, 24; SW, 25; PXA+, 25), AdamW (XL, 24; BN, 24), and Muon (CLL, 25; SW, 25; LLS, 25). In contrast, optimizers with cautious weight decay perform bilevel optimization, a framework from classical optimization Sol07a ; Sol07b ; SS (17) that has been recently explored in machine learning GLL (21); PMA (24).

### v2

@@ 1 Introduction
Optimization algorithms lie at the core of modern deep learning, shaping not only convergence speed but also training stability and generalization ability across domains such as natural language processing and computer vision. As models and datasets scale, traditional methods such as stochastic gradient descent (SGD) and SGD with momentum (Sutskever et al., 2013) encounter limitations, including slow convergence in non-convex landscapes, sensitivity to learning rate schedules, and poor robustness to sparse or noisy gradients (Scaman and Malherbe, 2020; Zhao et al., 2025). In response, a wide range of alternatives have emerged, including adaptive gradient methods (Duchi et al., 2011; Kingma and Ba, 2015), approximate second-order approaches (Martens and Grosse, 2015; Gupta et al., 2018; Yao et al., 2021; Liu et al., 2024; Nguyen et al., 2024; Wen et al., 2025), and specialized algorithms for extreme training regimes (Liang et al., 2024b; Luo et al., 2024; Xie et al., 2024; Huang et al., 2025; Zhang et al., 2025).
Among these advances, decoupled weight decay (Loshchilov and Hutter, 2019) has proven especially influential. In its general form, decoupled weight decay augments any optimizer update $\mathbf{u}_{t}$ with a decay term applied directly to the parameters, i.e.
@@
This technique improves training stability and generalization by preventing the adaptive learning rates from interfering with regularization, as exemplified by the success of AdamW in large model training (Brown et al., 2020; Dosovitskiy et al., 2021; Touvron et al., 2023) and the subsequent development of state-of-the-art optimizers such as Lion (Chen et al., 2023), Lion-$\mathcal{K}$ (Chen et al., 2024), and Muon (Jordan et al., 2024; Liu et al., 2025).
However, decoupled weight decay remains agnostic to the directional alignment between the optimizer update and the parameters, which may hurt performance when they conflict. Intuitively, when the update $\mathbf{u}_{t}$ and parameters $\mathbf{x}_{t}$ point in the same direction for a given dimension, weight decay acts as a regularizer that improves stability; however, when their directions differ, applying decay actively resists beneficial movement toward the optimum. Furthermore, decoupled weight decay has been shown to implicitly impose regularization terms on the objective function (Chen et al., 2024; Xie and Li, 2024), which corresponds to parameter norm constraints for AdamW, Lion, and Muon.
@@
$\bullet$ We use Lyapunov analysis to show that standard optimizers (SGD(M), Lion-$\mathcal{K}$, Adam) with cautious weight decay are asymptotically stable and unbiased, in the sense that they optimize the original loss rather than a regularized surrogate. The regularization effect of cautious weight decay instead becomes a bilevel objective of finding locally Pareto-optimal points within the stationary manifold (Figure 2). Furthermore, we show a convergence rate for discrete-time Adam with cautious weight decay in the smooth nonconvex setting under additional assumptions.
$\bullet$ In language modeling (OLMo et al., 2025; Kamath et al., 2025) and ImageNet classification (Deng et al., 2009), we observe that cautious weight decay generally accelerates convergence and lowers final validation loss for AdamW, Lion, and Muon (e.g., Figure 1). These improvements translate into higher zero-shot accuracy on standard benchmarks from 338M to 2B parameters and across architectures without retuning baseline settings ($\approx$20,000 NVIDIA H100 HBM3-80GB GPU hours for all experiments).
@@ 2.1 Decoupled weight decay
where $\beta_{1}$ and $\beta_{2}$ are momentum coefficients and $\nabla\mathcal{K}$ is a subgradient of $\mathcal{K}$. Examples include Lion when $\mathcal{K}=\left\lVert\cdot\right\rVert_{1}$ and Muon when $\mathcal{K}=\left\lVert\cdot\right\rVert_{\textup{tr}}$, where $\left\lVert\cdot\right\rVert_{\textup{tr}}$ denotes the nuclear norm when the parameters are treated as a matrix.
@@ 2.2 Implicit regularization effects of weight decay
This suggests that Lion-$\mathcal{K}$ optimizes the regularized objective $\widehat{f}$, an observation made by Chen et al. (2024). In the special cases of Lion and Muon, $\mathcal{K}^{*}$ is the 0-$\infty$ indicator function of a dual norm ball, corresponding to the constrained optimization problems
@@
A similar analysis for AdamW suggests that it solves the box-constrained problem of minimizing $f(\mathbf{x})$ such that $\left\lVert\mathbf{x}\right\rVert_{\infty}\leq\frac{1}{\lambda}$, but convergence cannot be established due to the lack of a Lyapunov function. For more discussion, see Appendix C and Xie and Li (2024).
@@ 3.1 Convergence to the stationary manifold
where $(\cdot)^{+}:=\max(0,\cdot)$. LaSalle’s invariance principle (LaSalle, 1960) states that the accumulation points of any trajectory lie within the union of trajectories $\mathbf{z}_{t}$ that satisfy $\tfrac{\textup{d}}{\textup{d}t}\mathcal{H}(\mathbf{z}_{t})=0$ for all $t\geq 0$. Consequently, we conclude that SGD with cautious weight decay produces trajectories that approach the stationary set $\{\mathbf{x}\mid\nabla f(\mathbf{x})=\mathbf{0}\}$ of the original loss. This holds because cautious weight decay is applied only in a secondary fashion and is automatically deactivated whenever it conflicts with the main objective, thereby ensuring that the loss landscape remains unbiased.
@@ 3.2 Sliding mode dynamics
Due to the discontinuity of $f_{t}$, the solution may not be well defined in the classical or Carathéodory sense, especially across switching surfaces. We therefore interpret solutions in the Filippov sense (Filippov, 1988), where a discontinuous ODE is formally a differential inclusion that specifies that $\dot{\mathbf{z}}_{t}$ belongs to the closed convex envelope of the discontinuous vector field, i.e.
@@
Figure 3: Toy objectives and trajectories. Left: $f(x,y)=((y-3)^{2}+(x-3)^{2}-1)^{2}$. Right: $f(x,y)=(y-3-(x-3)^{2})^{2}$. We compare Adam, AdamW, and Adam + CWD; AdamW and CWD use the same weight decay $\lambda$, and all other hyperparameters $(\eta,\beta_{1},\beta_{2},\epsilon)$ are identical. For both objectives, Adam converges to a generic point on the minimizer manifold, whereas AdamW converges to a solution of the box-constrained problem $\min_{x,y}f(x,y)$ subject to $\max\{x,y\}\leq\frac{1}{\lambda}$. In contrast, Adam + CWD converges to the Pareto front of the minimizer manifold.
3.3 Discrete-time analysis
Leveraging the Lyapunov functions in Table 1, we can extend our continuous-time analysis to obtain convergence guarantees for the discrete-time dynamics of various optimizers with cautious weight decay. As a concrete example, we provide in Appendix E an explicit convergence rate for discrete-time Adam with cautious weight decay.
@@ 4 Experiments
Overview. We evaluate CWD against three standard optimizers—AdamW, Lion, and Muon—on autoregressive language modeling and ImageNet classification. For Transformer models with similar architecture to Gemma (Kamath et al., 2025) with 338M, 986M, and 2B parameters in the Simply (Liang et al., 2025) codebase, we follow the Chinchilla compute-optimal scaling rule—20 tokens per parameter (TPP) (Hoffmann et al., 2022) and train on C4 (Raffel et al., 2020). For each size, we grid-search batch size, learning rate, weight decay, warmup ratio, and optimizer-specific hyperparameters for the baselines (AdamW, Lion, Muon); we then reuse the selected baseline settings for CWD without retuning (details in Appendix F). Under matched settings, CWD lowers final validation loss and improves zero-shot accuracy. On the OLMo codebase (OLMo et al., 2025), we further study an over-training regime—OLMo-1B trained on 100B tokens (100 TPP) from Dolma (Soldaini et al., 2024). Under matched settings, CWD lowers final validation loss and improves zero-shot accuracy (Table 4). We also observe similar gains on ImageNet (Deng et al., 2009) across ViT (Dosovitskiy et al., 2021) and ResNet (He et al., 2016).
@@
Ablations on masking. Table 2 tests whether the benefits arise from the amount of decay applied or from CWD’s structure. Replacing our mask with a time-matched Bernoulli “random mask” substantially degrades performance (e.g., $2.56\!\to\!2.82$ for AdamW, $2.42\!\to\!2.73$ for Muon), showing that simply reducing the frequency of decay is insufficient. Substituting the indicator with the gradient-based $\mathbb{I}(\mathbf{g}\mathbf{x}\geq\mathbf{0})$ also underperforms. Finally, $\lambda=0$ remains worse than tuned decay, illustrating that explicit regularization is helpful and CWD leverages it more effectively. We further verify the difference between CWD and SPD (Tian et al., 2024) with additional instruction-following fine-tuning experiments. We fine-tune on the Alpaca GPT-4 dataset (Peng et al., 2023), which contains 52,000 instruction–response pairs generated by GPT-4 (OpenAI, 2023), and evaluate on three benchmarks: MMLU (Hendrycks et al., 2021), comprising 57 tasks and 14,079 questions covering a broad range of world knowledge; AGIEval (Zhong et al., 2024), a human-centric benchmark with 9,316 instances targeting general reasoning and problem-solving skills; and WinoGrande (Sakaguchi et al., 2021), a large-scale commonsense reasoning dataset with 44,000 instances. We consider two base models, TinyLlama (Zhang et al., 2024) and Mistral-7B (Jiang et al., 2023), and compare LoRA (Hu et al., 2022), SPD (Tian et al., 2024), a layerwise “inner-product” variant of CWD using $\mathbb{I}(\left\langle\mathbf{u},\mathbf{x}\right\rangle\geq 0)$, and our proposed “elementwise” CWD. Table 5 reports accuracy on MMLU, AGIEval, and WinoGrande for TinyLlama and Mistral-7B.
@@
Model Method MMLU (5-shot) $\uparrow$ AGIEval (3-shot) $\uparrow$ WinoGrande (5-shot) $\uparrow$
TinyLlama LoRA (baseline) 25.81 $\pm$ 0.07 19.82 $\pm$ 0.11 61.33 $\pm$ 0.09
TinyLlama SPD 26.14 $\pm$ 0.08 20.21 $\pm$ 0.10 61.92 $\pm$ 0.08
TinyLlama Inner-product CWD 26.02 $\pm$ 0.08 19.80 $\pm$ 0.10 61.70 $\pm$ 0.09
TinyLlama Elementwise CWD 26.42 $\pm$ 0.09 20.12 $\pm$ 0.09 62.18 $\pm$ 0.08
Mistral-7B LoRA (baseline) 61.78 $\pm$ 0.09 27.56 $\pm$ 0.07 78.85 $\pm$ 0.11
Mistral-7B SPD 62.05 $\pm$ 0.08 27.98 $\pm$ 0.06 78.81 $\pm$ 0.10
Mistral-7B Inner-product CWD 61.76 $\pm$ 0.09 27.90 $\pm$ 0.07 78.83 $\pm$ 0.10
Mistral-7B Elementwise CWD 62.13 $\pm$ 0.07 28.31 $\pm$ 0.06 78.92 $\pm$ 0.09
Table 5: Accuracy on MMLU, AGIEval, and WinoGrande for TinyLlama and Mistral-7B fine-tuned on Alpaca GPT-4 using LoRA, SPD, and two variants of CWD. Both SPD and CWD improve over the LoRA baseline, and the proposed elementwise CWD matches or outperforms SPD and the inner-product variant on most benchmarks.
Scaling with model size. We measured final validation loss for AdamW and CWD (“Ours”) at  [^1] M, 338M, 986M, and 2B parameters on the same dataset (Figure 6). At every scale, CWD attains a lower final validation loss than AdamW, and the gap remains stable or even widens with model size, indicating that the advantages of cautious weight decay persist into the large-model regime.
Figure 6: Final validation loss versus model size for AdamW and Adam + CWD (“Ours”) on the same pretraining dataset, at  [^1] M, 338M, 986M, and 2B parameters. Across all scales, Adam + CWD achieves consistently lower final validation loss than AdamW, and the gap remains stable or slightly increases with model size, suggesting that the benefits of cautious weight decay persist in the large-model regime.
Model Size Optimizer Final Validation Loss
338M AdamW (baseline) 3.0136
338M Adam + CWD 3.0059
338M AdamC 3.0087
338M AdamC + CWD 2.9915
Table 6: Final validation loss for Gemma-based models with different optimizers at 338M parameters. Adding cautious weight decay improves both AdamW and AdamC, with AdamC + CWD achieving the lowest loss.
Model Size Optimizer Scheduler Final Validation Loss
338M AdamW (baseline) Cosine 3.0136
338M AdamW (baseline) WSD 3.0101
338M Adam + CWD Cosine 3.0059
338M Adam + CWD WSD 3.0014
Table 7: Final validation loss for a 338M model under different optimizer–scheduler combinations. Adam + CWD improves over the AdamW baseline for both cosine and WSD schedules, with the best performance obtained by Adam + CWD with WSD.
Compound improvement with other techniques. We also observe compounding gains when combining cautious weight decay with other optimizer techniques on Gemma-based model architectures. In particular, we compare AdamW, AdamC (Defazio, 2025), and their variants with CWD at 338M parameters (Table 6).
Robustness to learning-rate schedules. We observe that CWD improves performance for both cosine scheduling and the Warmup–Stable–Decay (WSD) schedule, which uses a 10% warmup, a long stable phase, and a 20% final decay (Table 7).
@@ 5 Related Work
Weight decay. Weight decay originated as an $\ell_{2}$ penalty for ill-posed problems and ridge regression (Tikhonov, 1963; Hoerl and Kennard, 1970) and was introduced to neural networks as a generalization tool to mitigate overfitting (Hanson and Pratt, 1988; Weigend et al., 1990; Krogh and Hertz, 1991). Loshchilov and Hutter (2019) showed that, for adaptive methods, weight decay and $\ell_{2}$ regularization are not equivalent, motivating the decoupled formulation in AdamW; subsequent work established decoupled decay as a standard feature of modern optimizers (Chen et al., 2023, 2024; Liu et al., 2025). Recent analyses suggest that in contemporary networks, weight decay functions more as a training accelerator and stabilizer than as explicit regularization (Krizhevsky et al., 2017; Hoffmann et al., 2022; Pan and Cao, 2023; D’Angelo et al., 2024). Interactions with early-stage training (Bjorck et al., 2021), normalization layers, and learning rate schedules (Defazio, 2025) have also been clarified, and architectural designs can obviate explicit decay (Loshchilov et al., 2025).
@@
Weight decay variants. Various efforts have been made to develop different adaptive variants of weight decay. For example, Xie et al. (2023) found that weight decay can lead to large gradient norms at the final phase of training and proposed Scheduled Weight Decay (SWD) to dynamically adjust weight decay strength based on gradient norms. Kosson et al. (2024) investigate how weight decay affects individual neuron updates, revealing rotational equilibrium states that balance learning across layers and neurons. Ghiasi et al. (2023) introduce adaptive weight decay that automatically tunes the hyperparameter during training based on classification and regularization loss gradients, achieving significant improvements in adversarial robustness. Tian et al. (2024) introduce Selective Projection Decay (SPD) for robust fine-tuning, featuring selective weight decay via a mask that is somewhat similar to CWD. However, SPD and CWD differ in significant ways, including the intended setting, mechanism, theoretical properties, and empirical performance.
@@
Masked or conditional updates. Several works have explored the sign-based conditioning of optimizer updates. Riedmiller and Braun (1993) introduced Rprop, which adjusted step sizes based on current gradient and past gradient sign agreement. Liang et al. (2024a) propose the cautious optimizer, which restricts updates to dimensions where the proposed update and current gradient share the same sign. Wang et al. (2024) apply a similar mask to Adam to improve robustness in online learning.
@@
Constrained and bilevel optimization. Decoupled weight decay can be interpreted through the lens of Frank–Wolfe algorithms for constrained optimization (Frank and Wolfe, 1956; Jaggi, 2013; Sfyraki and Wang, 2025; Pethick et al., 2025). This connection suggests that optimizers with decoupled weight decay implicitly solve constrained optimization problems, which was shown to be the case for Lion (Chen et al., 2024; Sfyraki and Wang, 2025; Pethick et al., 2025), AdamW (Xie and Li, 2024; Bernstein and Newhouse, 2024), and Muon (Chen et al., 2025; Sfyraki and Wang, 2025; Lau et al., 2025). In contrast, optimizers with cautious weight decay perform bilevel optimization, a framework from classical optimization (Solodov, 2007a, b; Sabach and Shtern, 2017) that has been recently explored in machine learning (Gong et al., 2021; Liu et al., 2022; Petrulionyte et al., 2024).